# Spectro-Photometry and Radial Distribution of Multiple Stellar Populations in Globular Clusters from Gaia XP Spectra

### v1

@@ Abstract
Understanding the formation of multiple populations in globular clusters (GCs) represents a challenge for stellar population studies. Nevertheless, the outermost GC regions, which may retain information of the initial configuration of the multiple populations, are poorly studied.
We use synthetic spectra that account for the chemical compositions of first- and second-population (1P, 2P) stars of 47 Tucanae to identify the spectral regions that are sensitive to its multiple populations. Hence, we defined new photometric bands that are efficient to disentangle 1P and 2P giant stars from Gaia XP spectra.
To test these new filters, we constructed the pseudo two-color diagrams dubbed chromosome maps (ChMs) and identified for the first time, 1P and 2P stars in the outermost cluster regions of 47 Tucanae and outside the tidal radius. We constructed similar diagrams for NGC 3201, NGC 6121, NGC 6752, and NGC 6397, thus exploring GCs with different metallicities. The ChMs allowed us to clearly disentangle 1P and 2P stars in the external regions of all clusters, with the exception of the metal-poor NGC 6397.
Our findings, together with literature results from more-internal regions, show that the 2P stars of 47 Tucanae and NGC 3201 are more-centrally concentrated than the 1P, whereas the multiple populations of NGC 6121, and NGC 6752 share the same radial distributions. These radial behaviors are consistent with the GC formation scenarios where 2P stars originate in the central regions. Noticeably, results on NGC 3201 are in tension with the conclusion from recent work that its 1P is more centrally concentrated than the 2P and might form with more central concentration.
@@ 1 Introduction
Globular clusters (GCs) with initial masses larger than $10^{5}M_{\odot}$ contain two main stellar groups. The first population (1P) exhibits a chemical composition comparable to that of field stars with similar metallicity. Conversely, the second population (2P) includes stars enhanced in He, N, and Na, and depleted in C and O (e.g. Kraft, 1994; Gratton et al., 2012; Bastian & Lardo, 2018; Milone & Marino, 2022). In addition to these light-element variations, some GCs (hereafter type II GCs) host stellar populations which differ in their heavy-element content (e.g. Yong & Grundahl, 2008; Marino et al., 2009; Marino et al., 2015; Milone et al., 2017; Johnson et al., 2015).
@@
Despite extensive observational and theoretical efforts, the mechanisms responsible for the formation of multiple stellar populations and the contributions of GCs to the early Galaxy’s assembly remain poorly understood. Various authors propose that the presence of multiple populations is linked to distinct star-formation events. In their scenarios, the 2P stars formed from material contaminated by more massive 1P stars, and the stars lost by the GCs provided a significant contribution to the Galactic halo (Dantona et al., 1983; D’Antona et al., 2016; Decressin et al., 2007; Denissenkov & Hartwick, 2014; Renzini et al., 2022; Lacchin et al., 2024). Alternatively, other scenarios suggest that 2P stars may arise either from binary-star evolution, or from accretion of polluted material onto already existing 1P stars. In these scenarios, GCs provided moderate contribution to the Galactic halo assembly (e.g. Wang et al., 2020; Gieles et al., 2018).
@@
The present-day radial distribution of 1P and 2P stars offers valuable insights into the origin and evolution of multiple populations. Specifically, GCs that have not fully undergone dynamic relaxation may preserve information regarding the initial configuration of 1P and 2P stars. Hence, it is crucial to study the multiple populations in the outermost cluster regions, which are characterized by long relaxation times.
Photometry proves to be an effective technique for exploring multiple stellar populations. The 1P and 2P stars define distinct sequences in appropriate photometric diagrams that are sensitive to their light-element abundances, thus providing the opportunity to study the multiple-population phenomenon in large stellar samples. Nowadays, photometric surveys of multiple populations have been performed by using Hubble Space Telescope (HST) observations of stars along various evolutionary phases, from the main-sequence (MS), to the red-giant (RGB), horizontal (HB) and asymptotic-giant giant (AGB) branch (e.g.  Milone et al., 2017, 2020; Lagioia et al., 2021; Dondoglio et al., 2021). The main limitation of these studies is that they are typically focused on the relatively small field of view of the HST cameras, and on the central cluster regions.
@@
Wide-field ground based photometry has been complementary to HST studies and it was instrumental to detect multiple stellar populations in the external GC regions (Marino et al., 2008; Yong et al., 2008; Milone et al., 2012a; Monelli et al., 2013; Jang et al., 2022; Lee, 2019, 2022). Nevertheless, these studies rarely approach the GC tidal radius and the cluster outskirts are nearly unexplored (but see Dondoglio et al., 2023, for recent photometric investigations of multiple populations outside the tidal radius of the GC NGC 1851).
@@
In this work, we focus on Type I GCs, which are characterized by stellar populations with similar metallicities but different contents of some light elements, including He, C, N, and O. The paper is organized as follows: Section 2 describes the data used in this paper. In Section 3, we use simulated spectra of 1P and 2P stars to define the most effective photometric bands to disentangle stellar populations with different light-element abundances in the Type I GC 47 Tucanae. These filters are used to construct photometric diagrams that allow us to identify 1P and 2P stars in GCs by using Gaia Data Release 3 low-resolution XP spectra. We first adopted 47 Tucanae as a test case, and then extended the analysis to the GCs NGC 6121, NGC 3201, NGC 6752, and NGC 6397. The radial distribution of the multiple stellar populations is discussed in Section 4, while Section 5 provides the summary and the discussion of the results.
@@ 2 Data
The main dataset that we used to investigate the multiple stellar populations in GCs consists in low-resolution spectroscopy, photometry, astrometry, and proper motions from Gaia data release 3 (DR3, Gaia Collaboration et al., 2023).
Gaia DR3 offers spectra with resolutions of about 30 to 100 for $\sim$200 million astronomical sources. Each spectrum, known as XP spectrum, comprises two distinct spectra obtained by the two components of the Gaia low resolution spectro-photometer. These spectra are referred to as ’blue photometer’ (BP) and ’red photometer’ (RP) spectra, covering the wavelength intervals of 3300–6800 Å and 6400–10500 Å, respectively.
@@
We used the stellar proper motions and parallaxes along with the parameters provided by Gaia DR3 that are indicative of the photometric and astrometric quality to select a sample of well-measured cluster members. To do this, we followed the recipe from our previous works (Milone et al., 2018; Cordoni et al., 2023). Moreover, we accounted for the contamination from nearby sources by using the BP/BR Excess Factor, $\epsilon$, which is a measure of the ratio between the sum of BP and RP, and G magnitudes and indicates the consistency of Gaia photometry. To select poorly-contaminated stars, the following constraint is applied (Gaia Collaboration et al., 2021):
@@
Photometry of cluster members in NGC 6121, NGC 3201, and NGC 6397 has been corrected for differential reddening through the approach outlined by Milone et al. (2012b, see their Section 2.1). In a nutshell, we derived the fiducial line of MS, sub-giant branch (SGB), and RGB stars in the $G_{\rm RP}$ vs. $G_{\rm BP}-G_{\rm RP}$ color-magnitude diagram (CMD) and calculated the color residuals from this fiducial. To assess the extent of differential reddening experienced by each star in the examined field of view, we assembled a sample of 35 neighbors consisting of bright MS, SGB, and faint-RGB cluster members without apparent binary features. Our most reliable estimate of differential reddening corresponds to the median of the color residuals, calculated along the reddening line. The reddening direction is derived by using appropriate absorption coefficients that we calculated as in Legnardi et al. (2023). We adopted the reddening laws corresponding to three values of the ratio, $R_{\rm V}$, between the absorption in the $V$ band, $A_{\rm V}$, and the difference between the absorption in the B and V bands, E(B--V). For NGC 6397, we assumed the canonical value, RV=3.1, whereas for NGC 6121 and NGC 3201 we adopted RV=3.85 and RV=3.19, respectively, as inferred by Legnardi et al. (2023) from multi-band HST photometry.
@@ 3 Hunting multiple populations in GCs with Gaia XP spectra
To investigate the multiple stellar populations in GCs, we defined a set of filters that maximise the separation between the sequences of 1P and 2P stars in the CMD. To do this, we first adopted in the Section 3.1 the GC 47 Tucanae as a test case, and then investigate in Section 3.2 three more metal-poor GCs, namely NGC 6121, NGC 6752, NGC 3201, and NGC 6397, thus exploring the efficiency of our method from [Fe/H]$\sim-0.75$ to [Fe/H]$\sim-2.03$ dex.
@@ 3.1 Multiple populations in 47 Tucanae
We used two synthetic spectra with [Fe/H]=--0.75 and [$\alpha$/Fe]=0.3 that represent two stars on the bottom of the RGB of 47 Tucanae. Both spectra have effective temperature Teff=4746 K, and gravity, $\log{g}=1.82$ but different abundances of C, N, and O. The first spectrum has solar carbon and nitrogen contents and [O/Fe]=0.3, similar with what is observed in 1P stars. The other spectrum represents a 2P star with extreme chemical composition and is enhanced in [N/Fe] by 1.2 dex and depleted in both [C/Fe] and [O/Fe] by 0.5 dex with respect to the 1P spectrum. The spectra are derived by Milone et al. (2023) by using the the SYNTHE programme (Kurucz & Avrett, 1981; Castelli, 2005; Kurucz, 2005) and the model atmospheres that they calculated with the ATLAS 12 computer program (Kurucz, 1970, 1993; Sbordone et al., 2004).
To illustrate the differences between the simulated spectra, we calculated the logarithm of the ratio between the fluxes of 2P and 1P spectrum multiplied by --2.5 and plotted the resulting quantity as a function of the wavelength, $\lambda$, in Figure 1. As expected, the most significant variations involve the regions around 3,400Å  and 3,800Å. Here, the 2P spectrum is more absorbed than the 1P one, owing to the stronger NH and CN molecular bands. In addition, around 4,300 Å  the 2P spectrum is less absorbed than the 1P one, due to the weaker CH bands.
To identify the photometric bands that are efficient to disentangle 1P and 2P stars, we defined a grid of magnitudes as:
@@
We calculated grids of magnitudes by varying the value of $\lambda_{0}$ from 3,200 to 10,000 Å  and by assuming four different values of $\delta\lambda=100,150,200,$ and 1,000 Å. For each combination of $\lambda_{0}$ and $\delta\lambda$, we calculated the magnitude difference between 1P and 2P stars. We found that the most-efficient filters to distinguish the distinct stellar populations of 47 Tucanae are F38000200 and F43000100, which are sensitive to N and C variations, respectively. In particular, the pseudo color $C_{\rm m3800,m4300,GRP}$=($m3800^{0200}-m4300^{0100}$)--($m4300^{0100}-G_{\rm RP}$) maximizes the separation between 1P and 2P stars. Indeed, the 2P stars, which are enhanced in N and depleted in C with respect to the 1P, exhibit smaller ($m4300^{0100}-G_{\rm RP}$) and larger ($m3800^{0200}-m4300^{0100}$) colors than 1P stars with the same luminosity. Magnitudes constructed with filters centered around $\lambda_{0}\sim 3,400$ AA, such as F34000100, would be very sensitive to stellar populations with different nitrogen abundances. Nevertheless, we excluded photometry in these bands from the analysis. Indeed, as already demonstrated by Cordoni et al. (2023), we verified that it is not possible to derive accurate determinations of the UV stellar magnitudes from the Gaia XP spectra, due to the poorer quality of the spectra in the extreme UV region.
@@
Figure 3: $\Delta_{{\it C}\rm m3800,m4300,GRP}$ vs. $\Delta_{\rm m4300,m9225}$ ChM for 47 Tucanae stars. The RGB stars are shown in the left panel, whereas the HB and AGB stars are plotted in the right panels with crimson and teal colors, respectively.
To better disentangle the multiple populations along the RGB of 47 Tucanae, we used the photometric diagrams of Figure 2 to construct the pseudo two-color diagram called chromosome map (ChM). In a nutshell, we derived the red and blue boundaries of the RGB and computed the $\Delta_{\rm{\it C}m3800,m4300,GRP}$ and $\Delta_{\rm m4300,m9225}$ pseudo colors by following the recipe by Milone et al. (2017, see their equations 1 and 2). The resulting $\Delta_{\rm{\it C}m3800,m4300,GRP}$ vs. $\Delta_{\rm m4300,m9225}$ ChM is plotted in the left panel of Figure 3. The color distribution in this ChM is not uniform. About half RGB stars are distributed around the origin of the ChM reference frame and have $\Delta_{{\it C}\rm m3800,m4300,GRP}$ values smaller than $\sim 0.2$ mag. The remaining RGB stars are distributed along a sequence that extends up towards large absolute values of $\Delta_{\rm{\it C}m3800,m4300,GRP}$ and $\Delta_{\rm m4300,m9225}$.
@@
Figure 4 compares the distribution of RGB stars in the ChM of Figure 3, which is reproduced in the top-left panel, and in other photometric and spectroscopic diagrams used in the literature to disentangle 1P and 2P stars in 47 Tucanae. The top-middle panel reproduces the $\Delta_{\rm{\it C}U,B,I}$ vs. $\Delta_{\rm B,I}$ ChM derived by Jang et al. (2022), where the 1P stars are clustered around the origin of this ChM, whereas 2P stars have $\Delta_{\rm{\it C}U,B,I}>0.1$ mag. The top-right panel shows $\Delta_{\rm{\it C}m3800,m4300,GRP}$ as a function of $\Delta_{\rm{\it C}U,B,I}$, thus combining photometry from our work and from Jang et al. (2022). In the bottom-left panel of Figure,4 we utilized the abundances of carbon and nitrogen relative to iron from APOGEE DR17 (Abdurro’uf et al., 2022) to illustrate the anticorrelation between [C/Fe] and [N/Fe]. In this diagram, the 1P stars exhibit nearly solar abundances of carbon and nitrogen, whereas 2P stars are depleted in carbon and enhanced in nitrogen.
@@
We used red dots and blue triangles to represent the stars with $\Delta_{\rm{\it C}\rm m3800,m4300,GRP}$ values smaller and larger than $\sim 0.2$ mag, respectively, for which either accurate $U,B,I$ photometry from (Jang et al., 2022) or spectroscopy from APOGEE DR17 is available. The fact that the bulk of the red stars, that we selected from the top-left ChM, populates the 1P sequence in the $\Delta_{\rm{\it C}U,B,I}$ vs. $\Delta_{\rm B,I}$ ChM demonstrates that they correspond to the first population of 47 Tucanae, whereas the selected blue stars mostly belong to the second population. This conclusion is corroborated by the correlation between the $\Delta_{\rm{\it C}m3800,m4300,GRP}$ and $\Delta_{\rm{\it C}U,B,I}$ (top-right panels) and by the fact that the bulk of the selected red stars share the same carbon and nitrogen abundances as 1P stars. In contrast, stars represented by blue triangles show depletion in carbon and enhancement in nitrogen, as expected for the 2P. We conclude that the $\Delta_{\rm{\it C}m3800,m4300,GRP}$ vs. $\Delta_{\rm m4300,m9225}$ ChM is an efficient tool to disentangle multiple populations with different light-element abundances in 47 Tucanae.
@@
Recent works have revealed an extended 1P sequence in the ChMs of 47 Tucanae derived from photometry collected either with the Hubble and James Webb telescopes or with ground-based facilities (Milone et al., 2017; Jang et al., 2022; Milone et al., 2023; Marino et al., 2024). The pseudo-color extension of 1P stars is due to metallicity variations among 1P stars of [Fe/H]$\sim 0.1$ dex (Legnardi et al., 2022; Marino et al., 2023). The lower-right panel of Figure 4 shows $\Delta_{\rm B,I}$ vs. $\Delta_{\rm m4300,m9225}$ for the probable 1P stars selected in the top-left panel alone. We defined two equally-populated groups of stars with different $\Delta_{\rm B,I}$ values, colored them aqua and orange, and represented the $\Delta_{\rm B,I}$ and $\Delta_{\rm m4300,m9225}$ histogram distribution of each group. The correlation between $\Delta_{\rm m4300,m9225}$ and $\Delta_{\rm B,I}$ and the evidence that the selected aqua and orange stars exhibit different average $\Delta_{\rm m4300,m9225}$ values corroborates the evidence that the color extension of the sequence is intrinsic, thus indicating the $\Delta_{\rm m4300,m9225}$ pseudo-color is an effective tool for detecting subtle metallicity variations among 1P stars.
3.2 Multiple populations in metal-poor GCs

### v2

@@ Abstract
Understanding the formation of multiple populations in globular clusters (GCs) represents a challenge for stellar population studies. Nevertheless, the outermost cluster regions, likely to hold clues about the initial configuration of GC stars, remain underexplored. We use synthetic spectra reflecting the chemical compositions of first- and second-population (1P, 2P) stars in 47 Tucanae to identify spectral regions sensitive to these populations. This led us to define new photometric bands that effectively distinguish 1P and 2P giant stars using Gaia XP spectra. Testing these filters, we constructed the pseudo two-color diagrams dubbed chromosome maps (ChMs) and, for the first time, identified 1P and 2P stars in the cluster’s outermost regions and beyond its tidal radius. We constructed similar diagrams for NGC 3201, NGC 6121, NGC 6752, and NGC 6397, thus exploring GCs with different metallicities. The ChMs effectively distinguished multiple populations in the outer regions of all clusters, except for the metal-poor NGC 6397. Our findings, together with literature results from more-internal regions, show that the 2P stars of 47 Tucanae are more-centrally concentrated than the 1P. A similar pattern is seen for 2P stars with extreme chemical composition of NGC 3201. The multiple populations of NGC 6121, and NGC 6752 share the same radial distributions. These radial behaviors are consistent with the GC formation scenarios where 2P stars originate in the central regions. Noticeably, results on NGC 3201 are in tension with the conclusion from recent work that its 1P is more centrally concentrated than the 2P and might form with more central concentration.
@@ 1 Introduction
Globular clusters (GCs) with initial masses larger than $\sim 10^{5}M_{\odot}$ contain two main stellar groups. The first population (1P) exhibits a chemical composition comparable to that of most Milky Way stars with comparable metal content. Conversely, the second population (2P) includes stars enhanced in He, N, and Na, and depleted in C and O (e.g. Kraft, 1994; Gratton et al., 2012; Bastian & Lardo, 2018; Milone & Marino, 2022). Moreover, some GCs (hereafter type II GCs) host stars which also differ in their heavy-element content (e.g. Yong & Grundahl, 2008; Marino et al., 2009; Marino et al., 2015; Milone et al., 2017; Johnson et al., 2015).
@@
Despite extensive observational and theoretical efforts, the mechanisms responsible for the formation of multiple stellar populations and the contributions of GCs to the early Galaxy’s assembly remain poorly understood. Various authors propose that the presence of multiple populations is linked to distinct star-formation events. In their scenarios, the 2P stars formed from material contaminated by more massive 1P stars, and the stars lost by the GCs provided a significant contribution to the Galactic halo (Dantona et al., 1983; D’Antona et al., 2016; Decressin et al., 2007; Denissenkov & Hartwick, 2014; Renzini et al., 2022; Lacchin et al., 2024). Alternatively, other authors suggest that 2P stars may arise either from binary-star evolution, or accumulating enriched material onto pre-existing 1P stars. In these scenarios, GCs provided moderate contribution to the Galactic halo assembly (e.g. Wang et al., 2020; Gieles et al., 2018).
@@
The current radial behavior of 1P and 2P stars offers valuable insights into the origins and evolution of multiple populations. Specifically, GCs that have not fully undergone dynamic relaxation may preserve information regarding the original distribution of these stellar populations (e.g. Vesperini et al., 2013, 2018, 2021; Hénault-Brunet et al., 2015; Dalessandro et al., 2019; Sollima, 2021). Hence, it is crucial to study the multiple populations in the outermost cluster regions, which are characterized by long relaxation times.
Photometry proves to be an effective technique for exploring multiple stellar populations. The 1P and 2P stars form distinct sequences in appropriate photometric diagrams that are sensitive to their light-element abundances, thus providing the opportunity to analyze these populations in large stellar samples.
Nowadays, photometric surveys of multiple populations have been performed by using Hubble Space Telescope (HST) observations of both main-sequence (MS) stars and of evolved stars belonging to the red-giant, horizontal, and asymptotic-giant branches (RGB, HB, AGB, e.g.  Milone et al., 2017, 2020; Lagioia et al., 2021; Dondoglio et al., 2021). These works, limited by the narrow field of view of HST, have primarily explored the central regions of the clusters.
@@
Wide-field ground based photometry has been complementary to HST studies since it is instrumental to detect multiple stellar populations in the external GC regions (Marino et al., 2008; Yong et al., 2008; Milone et al., 2012a; Monelli et al., 2013; Jang et al., 2022; Lee, 2019, 2022). Nevertheless, these studies rarely approach the GC tidal radius while the cluster outskirts are nearly unexplored (but see Dondoglio et al., 2023, for recent photometric investigations of multiple populations beyond the tidal radius of NGC 1851).
@@
In this work, we focus on Type I GCs, which are characterized by stellar populations with similar metallicities but different contents of some light elements, including He, C, N, and O. The structure of the manuscript is as follows: In Section 2, we present our dataset. Section 3 exploits simulated spectra of 1P and 2P stars to define the most effective photometric bands to untangle stars with different elemental abundances in 47 Tucanae. These filters are used to construct photometric diagrams that allow us to detect 1P and 2P stars in GCs based on low-resolution XP spectra from Gaia Data Release 3 (DR3, Gaia Collaboration et al., 2023). We first adopted 47 Tucanae as a test case, and then extended the analysis to the GCs NGC 6121, NGC 3201, NGC 6752, and NGC 6397. The radial distribution of the multiple stellar populations is discussed in Section 4, while Section 5 offers a summary and discussion.
@@ 2 Data
The primary dataset we relied on to explore the GCs consists in low-resolution photometry, spectroscopy, astrometry, and proper motions, all sourced from Gaia DR3.
Gaia DR3 offers spectra with resolutions of about 30 to 100 for $\sim$200 million astronomical sources. Each spectrum, known as XP spectrum, comprises two distinct spectra obtained by the two components of the Gaia low resolution spectro-photometer. These spectra, termed ’blue photometer’ (BP) and ’red photometer’ (RP) cover the spectral intervals of 3300–6800 Å and 6400–10500 Å, respectively.
@@
We used the stellar proper motions and parallaxes along with the parameters provided by Gaia DR3 that are indicative of the photometric and astrometric quality to identify a group of accurately measured cluster members. To do this, we followed the recipe from our previous works (Milone et al., 2018; Cordoni et al., 2023). Moreover, we accounted for the contamination from nearby sources by using the BP/BR Excess Factor, $\epsilon$, which is a measure of the ratio between the sum of BP and RP, and G magnitudes and indicates the consistency of Gaia photometry. To select poorly-contaminated stars, the following constraint is applied (Gaia Collaboration et al., 2021):
@@
Photometry of cluster members in NGC 6121, NGC 3201, and NGC 6397 has been corrected for differential reddening through the approach outlined by Milone et al. (2012b, see their Section 2.1). To summarize, we calculated the ridge line for MS, sub-giant branch (SGB), and RGB stars in the $G_{\rm RP}$ vs. $G_{\rm BP}-G_{\rm RP}$ color-magnitude diagram (CMD) and used it a reference to derive the color residuals. To assess the extent of differential reddening experienced by each star in the examined field of view, we assembled a sample of 35 neighbors consisting of bright MS, SGB, and faint-RGB cluster members without apparent binary features. Our most accurate estimate of differential reddening is determined by the median of the color residuals, which are computed along the direction of the reddening vector. The reddening direction is derived by using appropriate absorption coefficients that we calculated as in Legnardi et al. (2023). We adopted the reddening laws corresponding to three values of the ratio, $R_{\rm V}$, between the absorption in the $V$ band, $A_{\rm V}$, and the difference between the absorption in the B and V bands, E(B--V). For NGC 6397, we assumed the canonical value, RV=3.1, whereas for NGC 6121 and NGC 3201 we adopted RV=3.85 and RV=3.19, respectively, as inferred by Legnardi et al. (2023) from multi-band HST photometry. As an example, we show the differential-reddening map in the direction of NGC 3201 in Figure 1.
Figure 1: Differential-reddening map in the direction of NGC 3201 (bottom left). The x-axis and y-axis are aligned with the directions of right ascension and declination, respectively, with the coordinates, $\delta$X and $\delta$Y, given relative to the cluster center. The top-left and bottom-right panels illustrate the reddening variation, $\Delta$E(B--V), relative to the average reddening of the analyzed stars, plotted as a function of the $\delta$X and $\delta$Y coordinates for stars in eight slices of $\delta$Y and $\delta$X, respectively.
@@ 3 Hunting multiple populations in GCs with Gaia XP spectra
To detect the distinct stellar populations, we defined a set of filters that maximize the separation of the 1P and 2P star sequences in the GC CMDs. To do this, we first adopted 47 Tucanae as a test case (Section 3.1), and then investigate more metal-poor GCs, namely NGC 6121, NGC 6752, NGC 3201, and NGC 6397 (Section 3.2). This allowed an exploration of the efficiency of our method from [Fe/H]$\sim-0.75$ to [Fe/H]$\sim-2.02$ dex.
@@ 3.1 Multiple populations in 47 Tucanae
We used two synthetic spectra with [Fe/H]=--0.75 and [$\alpha$/Fe]=0.3 that represent two stars on the bottom of the RGB of 47 Tucanae. Both spectra have an effective temperature Teff=4746 K, and a gravity, $\log{g}=1.82$ but differ in their contents of some light elements. The first spectrum has [C/Fe]=0.0, [N/Fe]=0.0 and [O/Fe]=0.3, similar with what is observed in 1P stars. The other spectrum represents a 2P star with extreme chemical composition and is enhanced in [N/Fe] by 1.2 dex and depleted in both [C/Fe] and [O/Fe] by 0.5 dex with respect to the 1P spectrum. The spectra are derived by Milone et al. (2023) by using the the SYNTHE programme (Kurucz & Avrett, 1981; Castelli, 2005; Kurucz, 2005) and the model atmospheres that they calculated with the ATLAS 12 computer program (Kurucz, 1970, 1993; Sbordone et al., 2004).
To illustrate the differences between the simulated spectra, we show the 2P/1P flux ratio, in logarithmic scale, against the wavelength, $\lambda$, in Figure 2. As expected, the most significant variations involve the regions around 3,400Å  and 3,800Å. Here, the 2P spectrum is more absorbed than the 1P one, owing to the stronger NH and CN molecular bands. In addition, around 4,300 Å  the 2P spectrum is less absorbed than the 1P one, due to the weaker CH bands.
To find the bands that are effective in photometrically distinguishing between 1P and 2P stars we defined a grid of magnitudes as:
@@
We calculated grids of magnitudes by varying the value of $\lambda_{0}$ from 3,200 to 10,000 Å  and by assuming four different values of $\delta\lambda=100,150,200,$ and 1,000 Å. For each combination of $\lambda_{0}$ and $\delta\lambda$, we calculated the magnitude difference between 1P and 2P stars. We found that the most-efficient filters to distinguish the distinct stellar populations of 47 Tucanae are F38000200 and F43000100, which are sensitive to N and C variations, respectively. In particular, the pseudo color $C_{\rm m3800,m4300,GRP}$=($m3800^{0200}-m4300^{0100}$)--($m4300^{0100}-G_{\rm RP}$) maximizes the separation between 1P and 2P stars. Indeed, for a fixed luminosity, the 2P stars, which are N-rich and C-poor, exhibit smaller ($m4300^{0100}-G_{\rm RP}$) and larger ($m3800^{0200}-m4300^{0100}$) colors compared to the 1P stars. Magnitudes constructed with filters centered around $\lambda_{0}\sim 3,400$ AA, such as F34000100, would be very sensitive to stellar populations with different nitrogen abundances. Nevertheless, we excluded photometry in these bands from the analysis. Indeed, as already demonstrated by Cordoni et al. (2023), we verified that it is not possible to derive accurate determinations of the UV stellar magnitudes from the Gaia XP spectra, due to the poorer quality of the spectra in the extreme UV region.
@@
Figure 4: $\Delta_{{\it C}\rm m3800,m4300,GRP}$ vs. $\Delta_{\rm m4300,m9225}$ ChM for 47 Tucanae stars. The left panel displays the ChM of RGB stars, whereas the HB and AGB stars are represented in the right panel with crimson and teal colors, respectively.
To better disentangle the multiple populations in the RGB of 47 Tucanae, we exploited the photometric diagrams of Figure 3 to construct the pseudo two-color diagram also known as chromosome map (ChM). In a nutshell, we derived the red and blue boundaries of the RGB and computed the $\Delta_{\rm{\it C}m3800,m4300,GRP}$ and $\Delta_{\rm m4300,m9225}$ pseudo colors by following the recipe by Milone et al. (2017, see their equations 1 and 2). The resulting $\Delta_{\rm{\it C}m3800,m4300,GRP}$ vs. $\Delta_{\rm m4300,m9225}$ ChM is plotted in the left panel of Figure 4. The color distribution in this ChM is not uniform. About half RGB stars are grouped near the origin of the ChM reference frame and have $\Delta_{{\it C}\rm m3800,m4300,GRP}$ values smaller than $\sim 0.2$ mag. The remaining RGB stars are distributed along a sequence that extends up towards large absolute values of $\Delta_{\rm{\it C}m3800,m4300,GRP}$ and $\Delta_{\rm m4300,m9225}$.
@@
Figure 5 compares the distribution of RGB stars in the ChM of Figure 4, which is reproduced in the top-left panel, and in other photometric and spectroscopic diagrams used in the literature to disentangle 1P and 2P stars in 47 Tucanae. The top-middle panel reproduces the $\Delta_{\rm{\it C}U,B,I}$ vs. $\Delta_{\rm B,I}$ ChM derived by Jang et al. (2022), where the 1P stars are clustered around the origin of this ChM, whereas 2P stars have $\Delta_{\rm{\it C}U,B,I}>0.1$ mag. The top-right panel shows $\Delta_{\rm{\it C}m3800,m4300,GRP}$ as a function of $\Delta_{\rm{\it C}U,B,I}$, thus combining photometry from our work and from Jang et al. (2022). In the bottom-left panel of Figure 5 we utilized the abundances of carbon and nitrogen relative to iron from APOGEE DR17 (Abdurro’uf et al., 2022) to illustrate the anticorrelation between [C/Fe] and [N/Fe]. In this diagram, the 1P stars exhibit nearly solar abundances of carbon and nitrogen, whereas 2P stars are depleted in carbon and enhanced in nitrogen.
@@
We used red dots and blue triangles to represent the stars with $\Delta_{\rm{\it C}\rm m3800,m4300,GRP}$ values smaller and larger than $\sim 0.2$ mag, respectively, for which either accurate $U,B,I$ photometry from (Jang et al., 2022) or spectroscopy from APOGEE DR17 is available. The fact that the bulk of the red stars, that we selected from the top-left ChM, populates the 1P sequence in the $\Delta_{\rm{\it C}U,B,I}$ vs. $\Delta_{\rm B,I}$ ChM demonstrates that they correspond to the first population of 47 Tucanae, whereas the selected blue stars mostly belong to the second population. This conclusion is corroborated by the correlation between the $\Delta_{\rm{\it C}m3800,m4300,GRP}$ and $\Delta_{\rm{\it C}U,B,I}$ (top-right panels) and by the fact that the bulk of the selected red stars share the same carbon and nitrogen abundances as 1P stars. In contrast, stars represented by blue triangles show depletion in carbon and enhancement in nitrogen, as expected for the 2P. We conclude that the $\Delta_{\rm{\it C}m3800,m4300,GRP}$ vs. $\Delta_{\rm m4300,m9225}$ ChM is an effective instrument to disentangle multiple populations in 47 Tucanae.
@@
Recent works have revealed an extended 1P sequence in the ChMs of 47 Tucanae derived from photometry collected either with the Hubble and James Webb telescopes or with ground-based facilities (Milone et al., 2017; Jang et al., 2022; Milone et al., 2023; Marino et al., 2024). The pseudo-color extension of 1P stars is due to metallicity variations among 1P stars of [Fe/H]$\sim 0.1$ dex (Legnardi et al., 2022; Marino et al., 2023). The lower-right panel of Figure 5 shows $\Delta_{\rm B,I}$ vs. $\Delta_{\rm m4300,m9225}$ for the probable 1P stars selected in the top-left panel alone. We defined two equally-populated groups of stars with different $\Delta_{\rm B,I}$ values, colored them aqua and orange, and represented the $\Delta_{\rm B,I}$ and $\Delta_{\rm m4300,m9225}$ histogram distribution of each group. The correlation between $\Delta_{\rm m4300,m9225}$ and $\Delta_{\rm B,I}$ and the finding that aqua and orange stars exhibit different average $\Delta_{\rm m4300,m9225}$ values corroborates the evidence that the color extension of the sequence is intrinsic, thus indicating the $\Delta_{\rm m4300,m9225}$ pseudo-color is an effective tool for detecting subtle metallicity variations among 1P stars.
Previous studies of 1P and 2P stars in the central regions of 47 Tucanae and various Galactic GCs have utilized the $\Delta_{C\rm F275W,F336W,F438W}$ vs.,$\Delta_{\rm F275W,F814W}$ ChM (e.g., Milone et al., 2017; Lagioia et al., 2024). For completeness, the left panel of Figure 6 presents the $\Delta_{C\rm F336W,F438W,F814W}$ vs. $\Delta_{\rm F438W,F814W}$ ChM of 47 Tucanae, constructed with filters similar to the Johnson-Cousins U, B, and I bands. The 1P and 2P stars selected by Milone et al. (2017) from the $\Delta_{C\rm F275W,F336W,F438W}$ vs. $\Delta_{\rm F275W,F814W}$ ChM (right panel of Figure 6) clearly occupy distinct regions in both ChMs. This demonstrates that results from U, B, I photometry for 47 Tucanae are directly comparable with literature findings based on F275W, F336W, F438W, and F814W data.
Figure 6: Comparison between the 47 Tucanae ChM constructed with the F336W, F438W, and F814W filters (left) and the one derived from F275W, F336W, F438W, and F814W magnitudes by Milone et al. (2017). The dashed line separates the bulk of 1P and 2P stars, which are colored red and blue, respectively.
3.2 Multiple populations in low-metallicity GCs